1. The Ghanaian economy typical of most economies in the world consists of individuals and institutions that hold money for different purposes. Given the current fractional reserve banking system being operated in the country where commercial banks out of excess reserves :



a. Derive an expression that shows how money supply is determined in the Ghanaian economy. (10marks)



b. Find the money multiplier and briefly explain the effects of currency to deposit ratio (cr) required reserve ratio (rr) and excess reserves (er) on the money supply process. (5marks)



c. Determine the value of money supply in the economy if total currency in circulation is ¢600 million and a sum of ¢200 million is held the books of the Bank of Ghana reserves. Other relevant estimates include: a currency to deposit ratio of 0.2, required reserve ratio of 0.1 and excess reserve ratio of 0.1 (5marks)




1
Expert's answer
2021-11-09T10:46:26-0500

Solution:

a.). Expression:

M "\times" V = P "\times" Q

Where: M = Money supply

            V = Velocity

            P = Price level

           Q = Real expenditures       


MS = MD = "P\times \frac{Q}{V}"

 

b.). Money multiplier = 1/Required reserve ratio

The currency deposit ratio displays the amount of currency that individuals hold as a proportion of aggregate deposits. An increase in deposit rates will motivate depositors to deposit more, hence leading to a decrease in the cash to aggregate deposit ratio, which will lead to a rise in the money multiplier.

 

The increase or decrease in the required reserve ratio determines the amount of cash that banks hold in reserves and thus money supply in the economy. When the central bank decreases the reserve ratio, it lowers the amount of cash that banks are required to hold in reserves, allowing them to make more loans thus increasing the money supply in the economy and vice versa.

 

When banks hold excess reserves, they loan out less money, thereby reducing the money supply.

 

c.). The value of money supply = Currency in circulation + Checkable deposits

Derive checkable deposits:

Currency deposit ratio = Currency in circulation "\div" Checkable deposits

0.2 = 600"\div" CD

Checkable deposits = 600"\div" 0.2 = 3,000 million

The value of money supply = 600 + 3000 = 3,600 million